A birds flies 3.0m west and then flies 5.0m south.



a) Draw the bird resultant displacement (d).



b) |d|



c) Direction of the bird’s resultant

1
Expert's answer
2021-11-01T12:53:52-0400

a) The sketch:



b) Determine the magnitude of displacement:


"|d|=\sqrt{3^2+5^2}=5.83\text{ m}."


c) The direction:


"\theta=\arctan\frac{5}{3}=59°\text{ S of W}."